# Safety Subspaces are Not Linearly Distinct: A Fine-Tuning Case Study

### v2

@@ 1 Introduction
Large Language Models (LLMs) show strong performance across a wide range of general-purpose tasks (47, 1, 55, 53, 62, 54). To ensure these models behave responsibly and align with human values, they undergo an additional process of safety alignment. This alignment is typically achieved during the post-training stage, enabling models to improve response quality, and refuse harmful prompts over the pre-trained stage. Despite known jailbreak methods that can bypass safeguards, aligned models are generally considered significantly safer than their base versions (40, 59, 48).
However, this alignment is fragile. Since safety is encoded in the model’s weights, any modification, such as further fine-tuning (FT), can compromise it. This exposes a deeper attack surface beyond prompt engineering: an adversary could insert a small number of malicious samples into a training set to subvert alignment (63, 64, 4, 68). Recent work shows that even benign FT, low-rank adaptation, or pruning can degrade a model’s safety profile (32, 46, 16, 17, 58). A growing line of research seeks to leverage the learned safety adherence to design defenses against adversarial attacks and interpret alignment mechanisms(3, 19, 61, 11, 33, 26, 30).
@@
To construct defenses, prior works (33, 19) have typically derived safety subspaces using one of two approaches: weight updates from general alignment (aligned–base model deltas) or updates from targeted safety tuning (safety–base model deltas). Our goal is to comprehensively investigate these so-called “safety subspaces” in order to determine whether the information they contain is truly specific to safety. If so, we could separate unsafe information from the useful knowledge learned during FT through simple projections, thereby ensuring that our fine-tuned models are both robust to safety and high-performing.
@@
Across all experiments, we uncover a consistent and surprising result: no subspace, whether in weight or activation space, captures safety-specific behavior in isolation. Although certain subspaces, such as those derived from alignment or safety-specific updates, are impactful, they amplify both safe and useful behaviors alike, indicating that safety is deeply entangled with general learning. Similarly, activations from harmful and helpful prompts occupy overlapping regions of activation space, providing no evidence for distinct safety-related regions. Together, these findings reveal a fundamental limitation of linear subspace-based strategies. Since safe and harmful behaviors cannot be cleanly separated linearly, then projection- or filtering-based defenses are unlikely to suppress harmfulness without incurring comparable losses in utility. Our key contributions are as follows:
@@
  • We find that orthogonal projection intended to filter harmful updates leads to proportional losses in utility, suggesting no selective linear separation between safe and useful behavior (Section 4).
@@
  • Finally, we reaffirm this hypothesis in activation space, showing that harmful prompts do not activate distinct linear regions (Section 6).
  • Through multiple experiments across five open-source LLMs from the Llama and Qwen families, we establish that safety alignment is not linearly separable, revealing fundamental limitations of subspace-based defenses.
@@ 2 Preliminaries
Alignment training typically emphasizes behavioral properties such as harmlessness, helpfulness, and honesty. Empirical studies (40, 19, 11) suggest that the alignment update $\Delta_{\mathrm{A}}$ encodes directions in parameter space that are strongly correlated with these safety attributes. This stage is also the sole point in production model training where safety is explicitly introduced into the model. Our goal is to systematically control the extent to which the subsequent task-specific update $\Delta_{\mathrm{T}}^{j}$ interacts with these alignment directions.
@@
To construct the alignment and safety subspaces, each tensor in the updates $\Delta_{\mathrm{A}}/\Delta_{\mathrm{S}}$ is first reshaped into a matrix (flattened if needed) $V_{\mathrm{A}},V_{\mathrm{S}}\in\mathbb{R}^{M\times N}$. From here on, we use $V_{\mathrm{A/S}}$ to refer to both $V_{\mathrm{A}},V_{\mathrm{S}}$. We perform a thin singular value decomposition (SVD) of the form $V_{\mathrm{A/S}}=U\Sigma V^{\top}$, which reveals the principal directions of parameter change (13, 39), ranked by their contribution to the Frobenius norm. The top $k$ (Top-K) right singular vectors in $V$ are then selected to define the alignment/safety-specific subspace:
@@ 4 Can Harmful Subspaces Be Removed?
Having analyzed helpful and harmful updates in isolation, we now turn to a more realistic scenario: contaminated FT. This setting involves adding a small fraction of harmful examples to an otherwise benign dataset, producing updates that blend both signals. Prior work has shown that even limited contamination can erode safety, causing models to revert to unsafe behaviors (64, 32, 46, 4, 68). While earlier experiments identified expressive subspaces, we now ask the reverse question: can harmful components of an update be removed? We test whether filtering specific subspaces, particularly those aligned with the dominant directions of the alignment or safety-tuned matrix, can reduce harmfulness while preserving utility.
Figure 3: Parallel projection-based update schemes across varying SVD fractions. We report the energy-kept ratio for models fine-tuned on Full Useful, Full Harmful and Contaminated data; and utility and harmfulness for models fine-tuned on Contaminated.
@@ 5 Are Safety Weight Subspaces Distinct?
Figures 4 and 8 (Appendix E) show the pairwise overlap between the dominant subspaces (top-$k$ directions) of each update. All pairs exhibit greater overlap than random baselines, indicating shared structure. However, in Figure 4, the strongest overlap is between the useful and harmful updates, rather than between alignment and harmful updates, as one might expect if safety were a shared component. This is a key finding. If a safety subspace existed, it would likely appear in the shared directions between alignment and harmful updates (or between safety and harmful updates), which affect safety in opposite ways. The absence of such overlap suggests that no consistent, linear safety-specific subspace exists. For safety-tuned models (Figure 8), the strongest overlap occurs between the useful and safety-specific updates, an even more counterintuitive result. This implies that, in terms of subspace overlap, the useful update lies closer to the safety-specific update than the harmful update does. Strikingly, this overlap is much larger than that between the harmful and safety-specific updates, even though, semantically, one might expect the latter to be most similar.
Implications: Shared Subspaces Drive Behavior, Not Safety.
@@ 6 Do Safety Subspaces Exist In Activation Space?
Figure 5 shows MSO values across all pairs of prompt categories. Useful and harmful prompts consistently exhibit overlap above the random baseline, indicating activation of shared high-energy subspaces in activation space. Interestingly, the overlap between the two harmful prompt sets is not always greater than their overlap with useful prompts; in some cases, the useful–harmful overlap exceeds the harmful–harmful overlap. The degree of overlap also varies across model configurations. Some models show strong alignment even in the top subspaces, while others exhibit more gradual increases, with overlap becoming significant only at higher energy thresholds. This variability suggests that representational similarity depends more on model-specific factors than on the safety content of the prompts. Additional results on other models are provided in Figure 9 (Appendix F).
Implications: Shared Subspaces Drive Behavior, Not Safety.
@@ Appendix A Related Work
Large Language Models (LLMs) do not inherently follow instructions and often exhibit socially undesirable behaviors. To address this, various post-training methods, instruction-tuning and reinforcement learning from human feedback, are applied to align base LLMs with human values and improve their instruction-following capabilities (40, 50, 59, 48). However, studies have shown that fine-tuning these aligned models on harmful data can undo this alignment, restoring their original, socially unacceptable behaviors (63). This unalignment phenomenon has been demonstrated in both open-source models (64, 32) and proprietary models (46, 4, 68) via publicly available fine-tuning APIs, thereby exposing a new attack surface (24, 10, 29). Moreover, even fine-tuning on benign downstream tasks can degrade alignment (17, 16).
@@
Post-processing stage defenses adjust the fine-tuned model to restore alignment and preserve usefulness. One approach adds a safety vector, defined as the difference between aligned and unaligned weights, to the fine-tuned parameters to regain safe behavior (3). Another line of work projects the fine-tuning update onto the alignment vector when their similarity drops below a threshold, or selectively merges layers from the fine-tuned and aligned models under the same criterion to achieve a similar effect (19, 11). A third strategy removes parameters identified as harmful after fine-tuning to restore alignment (20). It has been shown that safety directions in attention-head activations can also be located and used for targeted intervention (71) to realign the fine-tuned model. Another method detects update parameters whose signs contradict the original alignment and removes them (61). Additional work restores/finds safety-critical neurons (65, 6), fuses aligned and fine-tuned models (66), or adds an optimized post-hoc perturbation to recover alignment (57).
@@ Appendix D Can Harmful Subspaces Be Removed?
Figure 7: Parallel projection-based update schemes across varying SVD fractions. We report the energy-kept ratio for models fine-tuned on Full Useful, Full Harmful and Contaminated data; and utility and harmfulness for models fine-tuned on Contaminated.
@@ Appendix E Are Safety Weight Subspaces Distinct?
Figure 8: Mode Subspace Overlap (MSO) at the 70-, 80-, and 85- percentile layers for pairwise comparisons of the dominant subspaces from Harmful fine-tuned (H), Safety-Tuned (S), and Base (B) models.
@@ Appendix F Do Safety Subspaces Exist in Activation Space?
Appendix G Dataset Details

### v3

@@ 1 Introduction
Large Language Models (LLMs) show strong performance across a wide range of general-purpose tasks (51, 1, 63, 60, 70, 61). To ensure these models behave responsibly and align with human values, they undergo an additional process of safety alignment. This alignment is typically achieved during the post-training stage, enabling models to improve response quality, and refuse harmful prompts over the pre-trained stage. Despite known jailbreak methods that can bypass safeguards, aligned models are generally considered significantly safer than their base versions (44, 67, 53).
However, this alignment is fragile. Since safety is encoded in the model’s weights, any modification, such as further fine-tuning (FT), can compromise it. This exposes a deeper attack surface beyond prompt engineering: an adversary could insert a small number of malicious samples into a training set to subvert alignment (71, 72, 4, 76). Recent work shows that even benign FT, low-rank adaptation, or pruning can degrade a model’s safety profile (36, 50, 16, 17, 66). A growing line of research seeks to leverage the learned safety adherence to design defenses against adversarial attacks and interpret alignment mechanisms(3, 20, 69, 11, 37, 28, 33).
@@
To construct defenses, prior works (37, 20) have typically derived safety subspaces using one of two approaches: weight updates from general alignment (aligned–base model deltas) or updates from targeted safety tuning (safety–base model deltas). Our goal is to comprehensively investigate these so-called “safety subspaces” in order to determine whether the information they contain is truly specific to safety. If so, we could separate unsafe information from the useful knowledge learned during FT through simple projections, thereby ensuring that our fine-tuned models are both robust to safety and high-performing.
@@
Across all experiments, we find no evidence that any linear subspace-whether in weight or activation space-captures safety-specific behavior in isolation. Although certain subspaces, such as those derived from the principle components of alignment or safety-specific updates, are impactful, they amplify both safe and useful behaviors alike, indicating that safety is deeply entangled with general learning. Similarly, activations from harmful and helpful prompts occupy overlapping regions of activation space, providing no evidence for distinct safety-related regions. Together, these findings reveal a fundamental limitation of linear subspace-based strategies. Since safe and harmful behaviors cannot be cleanly separated linearly, then projection- or filtering-based defenses are unlikely to suppress harmfulness without incurring comparable losses in utility. Our key contributions are as follows:
@@
  • Finally, we reaffirm this hypothesis in activation space, showing that harmful prompts do not activate distinct linear regions We observe that safety and harmful updates do not exhibit relatively higher within-task subspace overlap than cross-task comparisons, providing no evidence for a weight-space region that isolates safety from general learning signals (Section 6).
  • Across multiple experiments on five open-source LLMs from the Llama and Qwen families, we consistently observe patterns inconsistent with linear separability of safety alignment, pointing to inherent limitations in subspace-based defenses.
@@ 2 Preliminaries
Alignment training typically emphasizes behavioral properties such as harmlessness, helpfulness, and honesty. Empirical studies (44, 20, 11) suggest that the alignment update $\Delta_{\mathrm{A}}$ encodes directions in parameter space that are strongly correlated with these safety attributes. This stage is also the sole point in production model training where safety is explicitly introduced into the model. Our goal is to systematically control the extent to which the subsequent task-specific update $\Delta_{\mathrm{T}}^{j}$ interacts with these alignment directions.
@@
To construct the alignment and safety subspaces, each tensor in the updates $\Delta_{\mathrm{A}}/\Delta_{\mathrm{S}}$ is first reshaped into a matrix (flattened if needed) $V_{\mathrm{A}},V_{\mathrm{S}}\in\mathbb{R}^{M\times N}$. From here on, we use $V_{\mathrm{A/S}}$ to refer to both $V_{\mathrm{A}},V_{\mathrm{S}}$. We perform a thin singular value decomposition (SVD) of the form $V_{\mathrm{A/S}}=U\Sigma V^{\top}$, which reveals the principal directions of parameter change (13, 43), ranked by their contribution to the Frobenius norm. The top $k$ (Top-K) right singular vectors in $V$ are then selected to define the alignment/safety-specific subspace:
@@ 4 Can Harmful Subspaces Be Removed?
Having analyzed helpful and harmful updates in isolation, we now turn to a more realistic scenario: contaminated FT. This setting involves adding a small fraction of harmful examples to an otherwise benign dataset, producing updates that blend both signals. Prior work has shown that even limited contamination can erode safety, causing models to revert to unsafe behaviors (72, 36, 50, 4, 76). While earlier experiments identified expressive subspaces, we now ask the reverse question: can harmful components of an update be removed? We test whether filtering specific subspaces, particularly those aligned with the dominant directions of the alignment or safety-tuned matrix, can reduce harmfulness while preserving utility.
Figure 3: Orthogonal projection-based update schemes across varying SVD fractions. We report the energy-kept ratio for models fine-tuned on Full Useful, Full Harmful and Contaminated data; and utility and harmfulness for models fine-tuned on Contaminated.
@@
We test whether projecting individual layers can cleanly separate safety and learning directions in Appendix I. Specifically, we repeat the two weight-space analyses described above and observe that no layer provides such separation.
@@ 5 Are Safety Weight Subspaces Distinct?
Figures 4 and 8 (Appendix E) show the pairwise overlap between the dominant subspaces (top-$k$ directions) of each update.
We report per-layer results in Figures 10, 11, 12, 13, and 14 (Appendix G), all of which show results consistent with our observations.
All pairs exhibit greater overlap than random baselines, indicating shared structure. However, in Figure 4, the strongest overlap is between the useful and harmful updates, rather than between alignment and harmful updates, as one might expect if safety were a shared component. This is a key finding. If a safety subspace existed, it would likely appear in the shared directions between alignment and harmful updates (or between safety and harmful updates), which affect safety in opposite ways. The absence of such overlap suggests that no consistent, linear safety-specific subspace exists. For safety-tuned models (Figure 8), the strongest overlap occurs between the useful and safety-specific updates, an even more counterintuitive result. This implies that, in terms of subspace overlap, the useful update lies closer to the safety-specific update than the harmful update does. Strikingly, this overlap is much larger than that between the harmful and safety-specific updates, even though, semantically, one might expect the latter to be most similar.
Implications: Shared Weight Subspaces Drive Behavior, Not Safety.
@@
We repeat all of our weight-space analyses for a model safety-tuned using only Direct Preference Optimization (DPO) in Appendix H, to assess whether our findings remain consistent across different safety training strategies. We find that all our results transfer cleanly to this setting.
@@ 6 Do Safety Subspaces Exist In Activation Space?
Figure 5 shows MSO values across all pairs of prompt categories. Useful and harmful prompts consistently exhibit overlap above the random baseline, indicating activation of shared high-energy subspaces in activation space. Interestingly, the overlap between the two harmful prompt sets is not always greater than their overlap with useful prompts; in some cases, the useful–harmful overlap exceeds the harmful–harmful overlap. The degree of overlap also varies across model configurations. Some models show strong alignment even in the top subspaces, while others exhibit more gradual increases, with overlap becoming significant only at higher energy thresholds. This variability suggests that representational similarity depends more on model-specific factors than on the safety content of the prompts. Additional results on other models are provided in Figure 9 (Appendix F). We report additional analysis experiments in Appendix K.
Implications: Shared Activation Subspaces Drive Behavior, Not Safety.
@@ 7 Conclusion
Reproducibility Statement
We place a strong emphasis on reproducibility. To this end, we make our implementation publicly available at https://github.com/CERT-Lab/safety-subspaces. A complete description of the experimental setup is provided throughout the paper in Sections 3, 4, 5 and 6, with additional hyperparameter details in Appendix B. All experiments are conducted on widely used, publicly available benchmark datasets, which we summarize in Appendix M.
@@ Appendix A Related Work
Large Language Models (LLMs) do not inherently follow instructions and often exhibit socially undesirable behaviors. To address this, various post-training methods, instruction-tuning and reinforcement learning from human feedback, are applied to align base LLMs with human values and improve their instruction-following capabilities (44, 56, 67, 53). However, studies have shown that fine-tuning these aligned models on harmful data can undo this alignment, restoring their original, socially unacceptable behaviors (71). This unalignment phenomenon has been demonstrated in both open-source models (72, 36) and proprietary models (50, 4, 76) via publicly available fine-tuning APIs, thereby exposing a new attack surface (25, 10, 32). Moreover, even fine-tuning on benign downstream tasks can degrade alignment (17, 16).
@@
Post-processing stage defenses adjust the fine-tuned model to restore alignment and preserve usefulness. One approach adds a safety vector, defined as the difference between aligned and unaligned weights, to the fine-tuned parameters to regain safe behavior (3). Another line of work projects the fine-tuning update onto the alignment vector when their similarity drops below a threshold, or selectively merges layers from the fine-tuned and aligned models under the same criterion to achieve a similar effect (20, 11). A third strategy removes parameters identified as harmful after fine-tuning to restore alignment (21). It has been shown that safety directions in attention-head activations can also be located and used for targeted intervention (79) to realign the fine-tuned model. Another method detects update parameters whose signs contradict the original alignment and removes them (69). Additional work restores/finds safety-critical neurons (73, 6), fuses aligned and fine-tuned models (74), or adds an optimized post-hoc perturbation to recover alignment (65).
@@ Appendix D Can Harmful Subspaces Be Removed?
Figure 7: Orthogonal projection-based update schemes across varying SVD fractions. We report the energy-kept ratio for models fine-tuned on Full Useful, Full Harmful and Contaminated data; and utility and harmfulness for models fine-tuned on Contaminated.
@@ Appendix E Are Safety Weight Subspaces Distinct?
Figure 8: MSO at the 70-, 80-, and 85- percentile layers for pairwise comparisons of the dominant weight subspaces from Harmful fine-tuned (H), Safety-Tuned (S), and Base (B) models.
@@ Appendix F Do Safety Subspaces Exist in Activation Space?
Figure 9: MSO across layers in the 65–90% depth range for pairwise comparisons of activations from Useful (U) and multiple Harmful (H1, H2) prompt sets.
Appendix G Can Safety Weight Subspaces Be Distinct Layerwise?
To complete the analysis in Section 5, we report per-layer results in Figures 10, 11, 12, 13, and 14. We observe that the strongest overlap is never between the alignment and harmful updates, contrary to what one might expect if safety were represented as a shared component. This pattern is consistent across all layers and all models.
Figure 10: MSO at all layers for pairwise comparisons of the dominant weight subspaces from Harmful fine-tuned (H), Aligned (A), and Base (B) models (Qwen-2.5 1.5B).
Figure 11: MSO at all layers for pairwise comparisons of the dominant weight subspaces from Harmful fine-tuned (H), Aligned (A), and Base (B) models (Qwen-2.5 3B).
Figure 12: MSO at all layers for pairwise comparisons of the dominant weight subspaces from Harmful fine-tuned (H), Aligned (A), and Base (B) models (Qwen-2.5 7B).
Figure 13: MSO at all layers for pairwise comparisons of the dominant weight subspaces from Harmful fine-tuned (H), Aligned (A), and Base (B) models (Llama-3.2 1B).
Figure 14: MSO at all layers for pairwise comparisons of the dominant weight subspaces from Harmful fine-tuned (H), Aligned (A), and Base (B) models (Llama-2 7B).
Appendix H Effect of Safety-Tuning via Direct Preference Optimization
In this section, we extend our study to include a different alignment method, specifically the widely-used Direct Preference Optimization (DPO, (52)). This allows us to study how overlap varies under a different alignment method. In doing so, we can determine whether the entanglement we observe is a fundamental property of the model’s architecture or a consequence of the alignment procedure, and whether it persists across different forms of safety training.
We repeat the full weight-space analyses from Sections 3, 4, and 5 for a Llama-3.2 1B model that we safety-tuned using DPO on the PKU-SafeRLHF dataset (30). We train the model for 5k steps on a 10k-example subset of the dataset.
The results are shown in Figures 15, 16, and 17. We observe that all of our previous hypotheses and analyses hold consistently across these experiments. Together, these findings suggest that the underlying geometric behavior we identify is robust to the choice of alignment method.
Figure 15: Parallel projection-based update schemes across varying SVD fractions (Llama-3.2 1B). We report the energy-kept ratio for models fine-tuned on Full Useful and Full Harmful data, utility for models fine-tuned on Full Useful, and harmfulness for models fine-tuned on Full Harmful. Model alignment was performed using DPO.
Figure 16: Orthogonal projection-based update schemes across varying SVD fractions (Llama-3.2 1B). We report the energy-kept ratio for models fine-tuned on Full Useful, Full Harmful and Contaminated data; and utility and harmfulness for models fine-tuned on Contaminated. Model alignment was performed using DPO.
Figure 17: MSO at all layers for pairwise comparisons of the dominant subspaces from Harmful fine-tuned (H), Aligned (A), and Base (B) models (Llama-3.2 1B). Model alignment was performed using DPO.
Appendix I Do Safety and Learning Decouple at Any Layer?
In this section, we investigate whether a layer-specific strategy exists for isolating safety and learning subspaces. Specifically, we ask whether projecting (k) components from a single layer yields different behavior than projecting the same number of components across all layers.
We test this on Qwen2.5-1.5B by repeating the experiments from Sections 3 and 4, but restricting the projection to individual layers. We evaluate approximately every third layer in the network, resulting in 10 distinct layers.
The results, shown in Figures 18 and 19, reinforce our earlier hypothesis: no individual layer exhibits a clean separation between safety and learning directions. This further strengthens our claim that safety and task-learning signals are geometrically entangled throughout the model.
Figure 18: Parallel projection-based update schemes across varying SVD fractions (when applied to a specific layer only). We report the energy-kept ratio for models fine-tuned on Full Useful and Full Harmful data, utility for models fine-tuned on Full Useful, and harmfulness for models fine-tuned on Full Harmful.
Figure 19: Orthogonal projection-based update schemes across varying SVD fractions (when applied to a specific layer only). We report the energy-kept ratio for models fine-tuned on Full Useful, Full Harmful and Contaminated data; and utility and harmfulness for models fine-tuned on Contaminated.
Appendix J Robustness Checks: General Ability and Over-Refusal
J.1 Assessing General Ability Confounds via MMLU
We want to test whether the harmfulness metric is confounded with general ability (instruction following and coherence). More concretely, we aim to determine whether the results reported in our paper could simply be due to a loss in general capability. To evaluate this, we use MMLU (Massive Multitask Language Understanding, (18)), a widely used benchmark for general ability. We repeat the experiments from Sections 3 and 4 and report results in Figures 20 and 21 for Qwen-2.5 1.5B and 3B.
We observe that MMLU accuracy remains stable at approximately 61–62% and 64–65%, respectively for the two models, across all projection settings, indicating no degradation in general ability. This strengthens our findings and confirms that our analysis is scientifically robust rather than an artifact of reduced model capability.
J.2 Ruling Out Over-Refusal as a Confounding Factor
We test whether projection induces over-refusal, which would indicate that our safety results might be confounded by models incorrectly rejecting benign queries. To evaluate this, we measure refusal rates on benign prompts using XSTest (55), which ideally should not be refused. We repeat the experiments from Sections 3 and 4 and report results in Figures 20 and 21 for Qwen-2.5 1.5B and 3B.
Across all projection settings, refusal rates remain very low, around 3–4% and 5–6% respectively for the two models. These consistently low rates indicate that our safety metrics are robust and that the findings reported in the paper are not confounded by projection-induced over-refusal.
Figure 20: Parallel projection-based update schemes across varying SVD fractions. We report a general-ability metric (MMLU accuracy) for models fine-tuned on Full Useful, and refusal rates on benign prompts (XSTest) for models fine-tuned on Full Harmful.
Figure 21: Orthogonal projection-based update schemes across varying SVD fractions. We report a general-ability metric (MMLU accuracy) and refusal rates on benign prompts (XSTest) for models fine-tuned on Contaminated.
Appendix K Additional Activation Space Analyses
We do several experiments to complement the discussion in Section 6.
K.1 Activation Space Analysis Across All Layers
We report activation-space results layer by layer to fully capture layer-specific behavior. These results are shown for Qwen-2.5 1.5B, in Figures 22, 23, and 24. We observe that our hypothesis and findings hold consistently across all layers.
Figure 22: MSO across all layers for activations from Useful (Math) and Harmful (BeaverTails) prompt sets.
Figure 23: MSO across all layers for activations from Useful (Math) and Harmful (ToxiGen) prompt sets.
Figure 24: MSO across all layers for activations from different Harmful (BeaverTails, ToxiGen) prompt sets.
K.2 Impact of Using Early Token Windows
We show the effect of using earlier token windows in Figure 25 for Qwen-2.5 1.5B and 3B. We observe that our results and hypothesis hold true in this setting as well.
Figure 25: Average MSO (for early token windows) across layers in the 65–90% depth range for pairwise comparisons of activations from Useful (U) and multiple Harmful (H1, H2) prompt sets.
K.3 Impact of Using Refusal Prompts
We investigate the effect of using a refusal-oriented dataset (harmless prompts from Extended Refusal, (57)) as the counterpart to harmful prompts. Instead of relying on a generic “useful” dataset such as MATH, this allows us to test whether our findings depend on the choice of benign data. As shown for Qwen-2.5 1.5B in Figure 26, the results remain consistent under this substitution.
Figure 26: Average MSO across layers in the 65–90% depth range for activations from Refusal and Harmful prompt sets.
Appendix L Distinction from Activation Steering Methods
Our claims are compatible with the empirical success of activation steering methods. However, they target a stronger assumption that is not established by those works: namely, that there exists a relatively low-dimensional, stable “safety subspace” whose directions uniquely encode safety and can be edited in isolation, without affecting acquired utility.
Prior work shows that for a given model checkpoint and prompt distribution, one can reliably extract directions or small subspaces in the residual stream whose manipulation strongly affects refusal or a unique safety characteristic. Examples include single refusal directions (2), conditional activation steering (34), manifold steering for overthinking (26), and safety directions that become easier to find in higher-dimensional models (62). These results demonstrate local linear controllability: in a given model, at specific layers, there are directions along which safety behavior is highly sensitive.
What they do not by themselves establish is that these directions constitute a unique, model-level safety subspace that is (i) conceptually specific to “safety” rather than general behavior, and (ii) robust under fine-tuning or other parameter changes.
At the same time, a complementary body of work points toward a more entangled and multi-directional picture: safety and helpfulness often trade off along shared directions (62), and safety behavior can be mediated by structured sets of neurons or layers rather than a single feature (78, 6).
Taken together, these works suggest the following picture: activation steering can reliably surface high-leverage (but local) directions that influence refusal within a fixed model, but this does not imply that safety is represented by a stable linear subspace. Instead, the evidence is consistent with safety being encoded through distributed mechanisms that vary across tasks and checkpoints. Our claims target this stronger, subspace-level interpretation, rather than the more limited form of local controllability demonstrated by steering methods. This is particularly important because a growing body of work (e.g., SafeLoRA, SaLoRA, etc.) relies on this stronger assumption, despite there being no evidence that local controllability extends to global safety behavior.
Appendix M Dataset Details